# Self-Evolved Reward Learning for LLMs

### v1

@@ Appendix B Algorithm Summary
Algorithm 1 summarizes the entire SSRL-LLM method, including iterative self-evolved RM training and reinforcement learning for LLM policy optimization.
@@
Input: Initial RM, unlabeled data $D_{\text{unlabeled}}$, validation data $D_{\text{val}}$, thresholds $\tau_{\text{low}}$, $\tau_{\text{high}}$, $\tau_{\Delta}$, $\delta$, learning rate $\eta$
@@
       Evaluate $R_{\theta}$ on $D_{\text{val}}$ to determine the learning status $\mathcal{S}$ using predicted probability differences and thresholds $\tau_{\text{low}}$, $\tau_{\text{high}}$, and $\tau_{\Delta}$;
@@
      /* Step 2: Filter Data and Retrain the Reward Model */
       Filter $D_{\text{unlabeled}}$ based on the current learning status $\mathcal{S}$ to obtain $D_{\text{filtered}}$ using Eq. 4;
      Retrain $R_{\theta}$ using pairwise loss $\mathcal{L}_{\text{pair}}$ (Eq. 5) on $D_{\text{filtered}}$ with learning rate $\eta$;
/* Step 3: Train the LLM via Reinforcement Learning */
Train LLM $\pi_{\phi}$ via RL using $R_{\theta}$ as the reward function and update $\pi_{\phi}$ with modified PPO (Eq. 6);
@@ C.1 Dataset statistics
Dataset Num Task Feedback type
Stackoverflow 31,284,837 QA human
HH-RLHF 169,352 QA human
UrtralFeedback 63967 QA GPT4
Summarize 179000 summarize human
@@ C.3 Training Details
PPO training. For PPO training, we use a learning rate of 1.4e-5 and set the generate sample length to 256. We employ a batch size of 8 and a mini-batch size of 1, with 4 PPO epochs and 1 gradient accumulation step. The target KL divergence is set to 0.1 and initial KL coefficient is set to 0.2. To ensure a more robust training process, we normalize the range of reward values to -0.5 to 0.5.

### v2

@@ Appendix B Algorithm
Algorithm 1 summarizes the entire SER method, including iterative self-evolved RM training and reinforcement learning for LLM policy optimization.
@@
Input: Initial RM $R_{\theta}$, unlabeled data $D_{\text{unlabeled}}$, human-labeled data $D_{\text{labeled}}$, thresholds $\tau_{\text{low}}$, $\tau_{\text{high}}$, $\tau_{\Delta}$, $\delta$, learning rate $\eta$
@@
/* Step 0: Pretrain the Reward Model */
Pretrain $R_{\theta}$ on the human-labeled data $D_{\text{labeled}}$ using pairwise loss $\mathcal{L}_{\text{pair}}$;
@@
       Evaluate $R_{\theta}$ on $D_{\text{unlabeled}}$ to determine the learning status $\mathcal{S}$ using predicted probabilities and thresholds $\tau_{\text{low}}$, $\tau_{\text{high}}$, and $\tau_{\Delta}$;
@@
      /* Step 2: Filter Data Based on Learning Status */
       if $\mathcal{S}=\text{Status}_{1}$ then
             Filter samples where predicted probabilities $p_{j}$ satisfy $p_{j}>\tau_{\text{high}}$ (confidently good) or $p_{j}<\tau_{\text{low}}$ (confidently bad) to construct $D_{\text{filtered}}$;
       else if $\mathcal{S}=\text{Status}_{2}$ then
             Filter paired samples where the absolute difference in predicted probabilities satisfies $|p_{j}^{1}-p_{j}^{2}|>\delta$ to construct $D_{\text{filtered}}$;
      /* Step 3: Update and Retrain the Reward Model */
       Update $D_{\text{filtered}}\leftarrow D_{\text{filtered}}^{n}+D_{\text{filtered}}^{n-1}$, where $D_{\text{filtered}}^{n}$ is the newly filtered data and $D_{\text{filtered}}^{n-1}$ is the data from the previous iteration.;
       Retrain $R_{\theta}$ on the updated $D_{\text{filtered}}$ using pairwise loss $\mathcal{L}_{\text{pair}}$ with learning rate $\eta$;
/* Step 4: Train the LLM via Reinforcement Learning */
Train LLM $\pi_{\phi}$ using $R_{\theta}$ as the reward function and update $\pi_{\phi}$ with modified PPO (Eq. 6);
@@ C.1 Dataset statistics
Dataset Num Task Feedback type Response type
Stackoverflow 31,284,837 QA human human response
HH-RLHF 169,352 QA human LLM response
UltraFeedback 63967 QA GPT4 LLM response
Summarize 179000 summarize human LLM&human response
@@ C.3 Training Details
PPO training. For PPO training, we use a learning rate of 1.4e-5 and set the generate sample length to 256. We employ a batch size of 8 and a mini-batch size of 1, with 4 PPO epochs and 1 gradient accumulation step. The target KL divergence is set to 0.1 and initial KL coefficient is set to 0.2. To ensure a more robust training process, we normalize the range of reward values to -1 to 1.
Thresholds. The thresholds $\tau_{\text{high}},\tau_{\text{low}}$, and $\tau_{\Delta}$ were determined through extensive hyper-parameter tuning to balance precision and recall in the self-training process. Specifically, we experimented with the following values:
  • $\tau_{\text{high}}\in\{0.55,0.65,0.75\}$
  • $\tau_{\text{low}}\in\{0.45,0.35,0.25\}$
  • $\tau_{\Delta}\in\{0.3,0.4,0.5\}$
After evaluating the RM’s performance with these parameters, we selected $\tau_{\text{high}}=0.55,\tau_{\text{low}}=0.45$, and $\tau_{\Delta}=0.3$ as they provided the most consistent improvements in the RM’s ability to self-label effectively without introducing significant error amplification.
@@ D.4 GPT4 evaluation Prompt
Appendix E Cost and Performance of SER vs. Human Labeling
The cost of using the SER method is significantly lower than employing human labeling, while achieving comparable performance. According to the following calculation method, the cost of the SER method is more than 6X lower than that of using human labeling.
E.1 Cost Estimation
Human Labeling Cost. Google Cloud’s human annotation service  [^1] https://cloud.google.com/ai-platform/data-labeling/pricing charges approximately \$0.11 USD / 50 words for classification tasks at the time of writting. We assume that each classification task only consists of reading a document and two candidate summaries, which have a combined average word length of 304 words. We estimate the human labeling cost per example to be \$0.668 USD (304 words *\$0.11 / 50 words) (Lee et al., 2023). The detailed calculation can be found in Equation 19.
[…]
LLM Labeling Cost. We evaluated the use of GPT-4o in place of human labels for the SER method. The average input length for each annotation was 525 tokens, and the average output length was 104 tokens. For the GPT-4o model  [^2] https://platform.openai.com/docs/pricing, the input cost is \$0.0025 per 1,000 tokens, and the output cost is \$0.01 per 1,000 tokens. Each response was scored three times by GPT-4o to provide a preference pair. Based on these parameters, the labeling cost per example was calculated as follows:
[…]
Inference Cost. Based on estimations using GPU pricing from Amazon Cloud  [^3] https://aws.amazon.com/pricing ( A machine equipped with 8 A100 GPUs incurs a cost of 32.77 USD per hour ), the average inference cost per sample amounts to 1.338e-4 USD/sample. Based on our testing, we inferred 1,530 samples using a single A100 GPU, which required 3 minutes. The detailed calculation of inference cost can be found in Equation 21. Our estimated SER cost per sample is \$0.10054 USD ( here we provide an approximate estimation. Per sample cost = 0.67*0.15 + 3*1.338e-4 USD. In SER, 1 inference is conducted per iteration, resulting in a total of 3 additional inferences).
[…]
E.2 Performance and Cost Comparison
Table 4: Performance and cost comparison in HH-RLHF dataset. In addition to comparing the 15% human-labeled data with SER, we further explored the effectiveness of replacing human-labeled data with LLM-labeled data. Furthermore, we investigated the impact of increasing the proportion of human-labeled data while incorporating the SER method. Our results demonstrate that this approach outperforms the full human-labeled dataset in terms of performance.
Method Cost per Sample(USD) Accuracy(%)
Full Human 0.668 70.45
15% Human-Labeled +SER 0.100 68.56
15% LLM-Labeled + SER 0.002 67.64
60% Human-Labeled + SER 0.402 71.83
SER utilized only 15% of the human-labeled data, resulting in a significant reduction in data dependency. In training stage, our computational costs are comparable to those incurred when using the full dataset, with additional inference costs being introduced solely during step 1 Self-label with Reward Model. We compared the performance and cost of SER and human labeling, as shown in Table 4. We observe that:
Cost-Effectiveness of SER. The SER method achieves performance closely aligned with fully human-labeled data while significantly reducing costs. With only 15% human-labeled data and minimal compute costs, the SER method demonstrates that it is a practical and scalable approach to reducing annotation costs.
Trade-Off Between Annotation and Compute Costs. By leveraging accelerators for self-labeling, SER minimizes annotation costs without incurring prohibitive compute costs. At ( K = 4 ) loops, SER achieves a strong balance between performance and cost, suggesting it is a viable alternative to full human labeling for various tasks.
Scalability. Uner lower cost, the (60% human labeled + SER) outperform the “Full Human” accuracy. The SER method provides a cost-efficient framework for scaling reward models without significantly compromising performance, making it suitable for real-world applications where annotation budgets are constrained.
Appendix F Benchmark Evaluation
To investigate in greater detail the impact of SER on reward models and preference alignment, we evaluated both the reward models and the models after PPO across multiple benchmarks. Specifically, for the reward models, evaluations were conducted on RewardBench (Lambert et al., 2024). For the preference alignment models, assessments were carried out on MT-Bench (Zheng et al., 2023) and Arena-Hard (Li et al., 2024). Overall, we found that the evaluation results on the benchmarks were consistent with those obtained using LLM-as-a-judge, and that SER significantly enhanced model performance under data-limited conditions.
F.1 Reward Modeling Results in RewardBench
As shown in Table 5, we present results for the RewardBench evaluation of models trained on the UltraFeedback dataset. This dataset offers a comprehensive benchmark, allowing us to evaluate generalization across diverse tasks.
Table 5: The table shows the performance of reward models trained on the Ultrafeedback dataset, evaluated on RewardBench. Here, ’Avg.’ denotes the average score. As demonstrated, under the same amount of human-annotated data, SER significantly outperforms Loop0 and achieves performance comparable to that obtained using the full human-annotated dataset.
Model Method Avg. Chat Chat-Hard Safety Reasoning
Llama 3 8B Loop 0 59.1 70.7 44.1 52.2 69.7
SER 72.3 97.2 58.2 67.8 75.0
Full Data 75.9 95.5 58.5 73.9 65.4
Mistral 7B Loop 0 56.3 55.9 51.3 59.4 63.2
SER 72.0 85.5 57.1 64.5 62.0
Full Data 66.8 93.8 52.4 64.1 60.3
Llama 2 13B Loop 0 56.3 82.7 45.2 66.0 59.0
SER 70.5 92.4 52.7 66.0 71.8
Full Data 74.1 95.5 54.1 68.7 63.7
The results demonstrate that SER significantly enhances model performance compared to the baseline Loop 0, achieving results close to those obtained with the full human-annotated dataset. In tasks related to dialogue (Chat and Chat-Hard), which are central to the UltraFeedback dataset, SER achieves performance nearly identical to models trained on the full dataset.
F.2 PPO Model Results in MT-Bench and Arena-Hard
To further evaluate the downstream performance of models trained with SER, we provide results from MT-Bench and Arena-Hard, comparing SER to both the Full Dataset and SFT baselines, as shown in Table 6 and Table 7.
Table 6: The performance of PPO models trained on the HH-RLHF dataset in MT-Bench.
Model Comparison Win Tie Lose
Llama 3 8B SER vs Full Dataset 96 (30.0%) 110 (34.3%) 114 (35.6%)
SER vs SFT 116 (36.25%) 112 (35.0%) 92 (28.8%)
Mistral 7B SER vs Full Dataset 61 (19.0%) 214 (66.9%) 45 (14.1%)
SER vs SFT 73 (22.8%) 199 (62.2%) 48 (15.0%)
Table 7: The performance of PPO models trained on the HH-RLHF dataset in Arena-Hard.
Model Comparison Win Tie Lose
Llama 3 8B SER vs Full Dataset 62 (12.4%) 363 (72.6%) 75 (15.0%)
SER vs SFT 70 (14.0%) 379 (75.8%) 51 (10.2%)
Mistral 7B SER vs Full Dataset 34 (6.8%) 442 (88.4%) 24 (4.8%)
SER vs SFT 41 (8.2%) 435 (87.0%) 24 (4.8%)
The results highlight that SER achieves competitive performance relative to models trained on the full dataset and often surpasses SFT-trained models (because Arena-Hard features more challenging test questions, it yields a relatively higher proportion of ties; however, improvements can still be observed). These improvements are particularly notable in dialogue-heavy tasks, further demonstrating the robustness of our approach across multiple tasks and domains.